Logiweb bibliography of peano axioms

Up Help

[0] peano axioms

[1] base

The pyk compiler, version 0.grue.20050603 by Klaus Grue GRD-2005-06-03.UTC:15:58:02.061312 = MJD-53524.TAI:15:58:34.061312 = LGT-4624531114061312e-6